### v1

# Stochastic Deep Learning Surrogate Models for Uncertainty Propagation in Microstructure-Properties of Ceramic Aerogels

Abstract
Deep learning surrogate models have become pivotal in enabling model-driven materials discovery to achieve exceptional properties. However, ensuring the accuracy and reliability of predictions from these models, trained on limited and sparse material datasets remains a significant challenge. This study introduces an integrated deep learning framework for predicting the synthesis, microstructure, and mechanical properties of ceramic aerogels, leveraging physics-based models such as Lattice Boltzmann simulations for microstructure formation and stochastic finite element methods for mechanical property calculations. To address the computational demands of repeated physics-based simulations required for experimental calibration and material design, a linked surrogate model is developed, leveraging Convolutional Neural Networks (CNNs) for stochastic microstructure generation and microstructure-to-mechanical property mapping. To overcome challenges associated with limited training datasets from expensive physical modeling, CNN training is formulated within a Bayesian inference framework, enabling robust uncertainty quantification in predictions. Numerical results highlight the strengths and limitations of the linked surrogate framework, demonstrating its effectiveness in predicting properties of aerogels with pore sizes and morphologies similar to the training data (in-distribution) and its ability to interpolate to new microstructural features between training data (out-of-distribution).
keywords:
Bayesian inference , convolutional neural networks , surrogate modeling , uncertainty quantification , ceramic aerogel
journal: Computational Materials Science
[…]
[inst1]organization=Department of Mechanical and Aerospace Engineering,
University at Buffalo,city=Buffalo, state=NY, country=USA
[…]
[inst2]organization=Computational and Data-Enabled Sciences,
University at Buffalo,city=Buffalo, state=NY, country=USA
1 Introduction
Ceramic aerogels, composed of interconnected amorphous ceramic nanoparticles, exhibit a highly porous structure and exceptional thermal insulation properties, making them ideal for advanced thermal management in extreme environments. Their applications span civil and military aviation [1], space exploration [2], fire-resistant wearables [3], microelectronics [4, 5], and energy-efficient building systems [6, 7]. However, widespread adoption is hindered by their low mechanical stability. Overcoming this requires model-driven materials discovery [8, 9, 10] to predict synthesis-microstructure-property relationships for optimizing synthesis parameters to create mechanically robust aerogels without compromising insulation performance.
Addressing computational challenges in uncertainty quantification, inference, and optimization for model-driven materials discovery hinges on developing surrogate models that approximate synthesis-to-property relationships. Trained on high-fidelity data, these models enable faster and more efficient predictions. Convolutional Neural Networks (CNNs) have become prominent in computational materials science [11, 12, 13, 14, 15], in generating realistic microstructural images and predicting macroscopic properties. However, their deployment in high-consequence applications, such as materials discovery, faces challenges, including sparse, limited, and stochastic training datasets from physical simulations. Traditional training methods relying on maximum likelihood estimation are prone to overfitting, producing overconfident and unreliable predictions. Furthermore, current approaches for ensuring CNN reliability depend on large datasets, limiting their effectiveness in small-data regimes. Framing training as a Bayesian inference problem offers a robust solution to handling small and uncertain datasets in deep learning surrogate modeling for physical simulations [16, 17]. This approach represents parameter uncertainty through probability distributions, reducing overfitting and enabling uncertainty quantification [18, 19]. However, applying Bayesian inference to CNNs is computationally challenging due to the large number of weight parameters. Zhu et al. [20] addressed this by employing a variational gradient descent algorithm based on Stein’s method, extending Bayesian inference to convolutional encoder-decoder networks. Their approach captured uncertainty in flow through heterogeneous media, even in scenarios where the input (permeability) and output (flow and pressure) fields lacked shared underlying structures. Similarly, Shridhar et al. [21] applied variational inference to estimate posterior weight distributions in CNNs, effectively propagating uncertainty in classification tasks across datasets like MNIST and CIFAR-100.
This work introduces an integrated modeling framework (Figure 1) for simulating the synthesis, microstructure, and properties of ceramic aerogels with high computational efficiency. Specifically, it focuses on a novel synthesis method utilizing in situ bubble-supported pore formation with foaming agents [22, 3], which addresses the limitations of conventional supercritical drying by preserving the porous structure without collapsing the solid matrix. This synthesis enables ambient-pressure, room-temperature drying, significantly reducing processing time while facilitating scalable and cost-effective production of ceramic aerogels. The synthesis-microstructure physics-based model in Figure 1, relies on Lattice Boltzmann simulations of the foaming process, capturing stochastic bubble nucleation, bubble dynamics, and coalescence to predict microstructure formation. The microstructure-property model employs stochastic finite element simulations to analyze the elastic deformation of the solid phase of the aerogel. Using the microstructural patterns, this model computes the displacement and strain fields to estimate the mechanical properties of the aerogel, i.e., the strain energy as the quantity of interest (QoI) for the modeling framework. To enable efficient simulations of synthesis, microstructure, and properties, we develop linked surrogates that demonstrate superior representations of the coupled system compared to composite surrogate models directly mapping synthesis inputs to QoI [23]. The microstructure surrogate leverages a scalable generative network trained on microscopic images, enabling the generation of morphologically consistent microstructures across larger domains necessary for macroscopic property predictions. Additionally, the surrogate model of the material property employs a CNN to map microstructural images to the QoI. A key innovation of this framework is the formulation of CNN training as Bayesian inference, enabling the estimation of the probability distribution of weight parameters using high-fidelity training data. To address the computational challenges of Bayesian inference in high-dimensional parameter spaces, we extend the Variational Inference approach used in Bayesian Neural Networks to accommodate the complexity of CNN architectures. The resulting Bayesian Convolutional Neural Network (BayesCNN) mitigates overconfidence in parameter estimation, supports training on small and uncertain datasets derived from high-fidelity simulations, and quantifies the uncertainty in surrogate model predictions. The linked surrogates are systematically employed to characterize and propagate various sources of uncertainty through the synthesis-microstructure-property relationship and the computational prediction of the QoI. This includes capturing microstructural randomness and quantifying errors introduced during surrogate model construction.
Figure 1: Flowchart of the integrated modeling framework developed in this work for propagating uncertainty through synthesis inputs $\boldsymbol{\xi}$, microstructure spatial field $\varphi(\mathbf{x})$, and mechanical property $Q$ of ceramic aerogels. The physics-based models, $\mathcal{F}_{LB}(\cdot)$ and $\mathcal{F}_{FE}(\cdot)$, are parameterized by $\boldsymbol{\theta}_{LB}$ and $\boldsymbol{\theta}_{FE}$, and utilize synthesis inputs and microstructural patterns to predict material properties. Computational efficiency is achieved through surrogate models $\mathcal{G}_{GAN}(\cdot)$ and $\mathcal{G}_{CNN}(\cdot)$, constructed using high-fidelity training data from physical simulations. The parameter of the surrogate models include the stochastic latent vector $\boldsymbol{z}$, deterministic weights of the generative network $\boldsymbol{w}_{GAN}$, and posterior probability distributions of the wights $\boldsymbol{w}_{CNN}$ from the Bayesian inference.
The remainder of this manuscript is structured as follows. Section 2 provides an overview of the ceramic aerogel synthesis process, including its modeling using Lattice Boltzmann simulations and the development of surrogate models to represent microstructural patterns. Section 3 describes the stochastic finite element model used to establish microstructure-property relationships and presents the comprehensive formulation and efficient solution algorithm of BayesCNN for surrogate modeling. Section 4 outlines the construction of individual surrogate models, detailing the creation of training datasets from high-fidelity simulations and the validation of these models. Additionally, this section demonstrates the application of the linked-surrogates for propagating of uncertainty from synthesis parameters and stochastic microstructure to the QoI. Finally, Section 5 presents a discussion of the findings and the concluding remarks.
2 Physical and surrogate models of aerogel microstructure
The exceptional thermal insulation of silica aerogels arises from their high porosity and nanoscale pore structure, which limits gas molecule movement. However, widespread use in engineering applications is hindered by the high production costs of current methods, such as supercritical or freeze-drying, e.g., [24]. To overcome these limitations, a novel synthesis process has been developed, utilizing in-situ bubble-supported pore formation for ambient pressure, room-temperature drying [22, 3, 25]. This method employs a silica precursor solution with cetrimonium bromide (CTAB) and an aqueous urea-based foaming agent. Hydrolysis of tetraethyl orthosilicate (TEOS) at CTAB micelle interfaces, combined with urea decomposition and ammonia and carbon dioxide release, generates in-situ bubbles, producing aerogels with randomly distributed pore sizes and morphologies. The primary synthesis inputs $\boldsymbol{\xi}$ include the concentration of TEOS ($C_{TEOS}$), which regulates the solid phase volume fraction and thereby affects the porosity and pore size of the aerogel. The concentration of CTAB ($C_{CTAB}$) plays a critical role in bubble formation, influencing the resulting pore size distribution. The concentration of urea ($C_{urea}$), acting as a foaming agent, reduces surface tension between the precursor solution and gaseous bubbles, thereby impacting the pore distribution. Additionally, the concentration of viscosifier ($C_{vis}$) determines the stability and morphology of the pore structure during the foaming process. These parameters collectively control the microstructural characteristics and properties of the ceramic aerogel.
2.1 Lattice Boltzmann simulation of microstructure formation
The computational modeling of the ceramic aerogel synthesis process, as discussed in the previous section, requires the integration of key physical phenomena involved in pre-aerogel foaming. These include bubble nucleation, gas diffusion into bubbles, bubble dynamics and coalescence, and the rheological behavior of the precursor, which often exhibits non-Newtonian characteristics. In this study, we utilize the Free Surface Lattice Boltzmann Method (FSLBM) for foaming simulations, e.g., [26, 27, 28, 29]. Specifically, we employ the open-source software package LBfoam [30], which is implemented using the Parallel Lattice Boltzmann (Palabos) library [31] for fluid dynamics and integrates a Volume-of-Fluid technique for accurate interface tracking. A brief overview of the fundamental equations is provided here for context, while detailed equations and solution algorithms can be found in [30].
The core of FSLBM is the Lattice Boltzmann simulation of complex fluid systems, e.g., [32, 33]. the computational domain is discretized into a regular lattice, where the particle distribution function $f_{i}(\mathbf{x},t)$ represents the expected number of particles at lattice site $\mathbf{x}$ and time $t$ in the state $i=0,1,\cdots,N$ defined by a set of discrete velocities $\mathbf{v}_{i}$. The discretized Lattice Boltzmann equation is
$f_{i}(\mathbf{x}+\mathbf{v}_{i}\Delta t,t+\Delta t)=f_{i}(\mathbf{x},t)+B(\mathbf{x},t),$ (1)
with the Bhatnagar-Gross-Krook (BGK) collision operator [33] as,
$B(\mathbf{x},t)=-\frac{\Delta t}{\tau}\left(f_{i}(\mathbf{x},t)-f_{i}^{eq}(\mathbf{x},t)\right),$ (2)
where $\Delta t$ is the time step, $\tau$ is a relaxation time, and $f_{i}^{eq}$ is the distribution function at equilibrium [33]. The solution of Eq. (1) involves two iterative steps: streaming, where the distribution functions $f_{i}$ move to neighboring lattice sites, and collision, where $f_{i}$ is locally updated at each site. Macroscopic fluid properties, such as density and velocity, are then derived from the moments of $f_{i}$. Building on this foundation, FSLBM simulations utilize a fast mass-tracking algorithm [27] to monitor the interfaces between gas and liquid phases throughout the computational domain. The liquid-gas interface is modeled as a free surface, with the VOF method employed to track the interface using a level function, which represents the liquid fraction within each computational cell. Surface tension $\sigma_{LB}$ are incorporated by calculating the interface curvature from the level function and applying the resulting pressure jump across the interface. In this work, to capture the highly viscous and shear-thinning behavior of the aerogel precursor, governed by the concentration of viscosifier $C_{vis}$, the Carreau-Yasuda model [34] is adopted to describe the fluid’s rheological properties. The model provides an accurate representation of the non-Newtonian characteristics essential for simulating the foaming process. The influence of TEOS concentration on bubble growth is incorporated by modeling gas diffusion into the bubbles and within the liquid phase using the advection-diffusion equation for gas concentration $C$,
$\frac{\partial C}{\partial t}+\mathbf{v}\cdot\nabla C=D\nabla^{2}C+S,$ (3)
where $\nabla$ represents the spatial gradient operator, $D$ denotes the diffusion coefficient, directly linked to $C_{TEOS}$, and $S$ is a source term representing gas generation or absorption as a function of $C_{urea}$, which regulates the rate of bubble growth.
Figure 2 presents snapshots of pre-aerogel foaming simulations conducted using FSLBM, illustrating the effects of varying foaming agent concentrations on the resulting ceramic aerogel microstructures. The simulations were initialized with a fluid pool of height 400 $\mu m$, with bubbles of three distinct nuclei sizes (8 $\mu m$, 18 $\mu m$, and 25 $\mu m$) distributed within the simulation domain using a Poisson disk-sampling algorithm [35]. For all simulations, the fluid density was set to 1 $kg/{\mu m}^{3}$, ambient pressure to 0.33 $N/{\mu m}^{2}$, surface tension to $\sigma_{LB}=5\times 10^{-3}$ $N/\mu m$, and the advection-diffusion source term to $S=5\times 10^{-5}$ ${\mu m}^{2}/s$. The variation in the number of initial bubbles and their radii led to distinct classes of ceramic microstructures with different average pore size at the final simulation time. As demonstrated later, the use of the Poisson disk-sampling algorithm enables FSLBM to generate multiple ensemble realizations for each microstructure class with the same numbers and initial radius of the bubbles to account for stochasticity of the aerogel microstructure.
Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Figure 2: FSLBM simulations of the ceramic aerogel foaming process (pores in white) showing the time evolution of bubbles’ growth and coalescence in the pre-aerogel to form a porous microstructure. All simulations are conducted over 20 minutes and 5000 timesteps each with $\Delta t=2$s and all the physical parameters are kept unchanged except the initial number of bubbles $n_{bub}$ and initial bubble radius $r_{bub}$ leading to different average aerogel pore size at the end of simulation: (a) $n_{bub}=250$, $r_{bub}=8$, leading to pore radius $8\mu m$; (b) $n_{bub}=50$, $r_{bub}=18$, leading to pore radius $18\mu m$; (c) $n_{bub}=25$, $r_{bub}=25$, leading to pore radius $25\mu m$. The simulation domain height was 550 $\mu$m and width was 400 $\mu$m, with 400 $\mu$m fluid pool height throughout the simulations.
2.2 Generative surrogate model of aerogel microstructure
FSLBM simulations of aerogel microstructures are computationally intensive, making tasks such as uncertainty quantification, requiring numerous evaluations of $\mathcal{F}_{LB}(\boldsymbol{\xi};\boldsymbol{\theta}_{LB})$ prohibitively expensive. To enable computationally efficient synthesis-microstructure mapping, we propose a surrogate model for FSLBM simulations based on the Generative Adversarial Network (GAN) introduced by [36]. GANs employ two deep convolutional neural networks: the generator $\mathcal{G}_{GAN}(\cdot)$ and the discriminator $\mathcal{D}(\cdot)$. These networks are trained an adversarial manner in a two-player minimax game, where the generator, parameterized by $\boldsymbol{w}_{GAN}$, maps a random latent input vector to the data space $p_{G}$, approximating the real data distribution $p_{data}$. Meanwhile, the discriminator, with weight parameters $\boldsymbol{w}_{dis}$, acts as a classifier, outputting the probability of a sample being real or generated. In this unsupervised learning setup, the generator minimizes the divergence between $p_{G}$ and $p_{data}$, while the discriminator maximizes it, until both networks reach a Nash equilibrium. Once trained on microstructural imaging data, the generator can efficiently produce ensembles of microstructures that closely replicate the characteristics of the training data. In recent years, the application of GANs to imaging and digital material data has gained significant attention, demonstrating their potential in capturing realistic microstructural features, e.g., [37, 38, 39, 40, 41].
To formulate the generative surrogate model for ceramic aerogel microstructures, we define a microstructure indicator function $\varphi(\mathbf{x})$, as illustrated in Figure 1. This spatial field assigns values of $\varphi=0$ to the spatial points $\mathbf{x}$ within the pores and $\varphi=1$ to those within the solid skeleton of the aerogel. The objective of the surrogate model is to generate a large ensemble of aerogel microstructural patterns, such that
$\varphi(\mathbf{x})=\mathcal{G}_{GAN}(\boldsymbol{\xi},\boldsymbol{z};\boldsymbol{w}_{GAN}),$ (4)
where $\mathcal{G}_{GAN}(\boldsymbol{\xi},\boldsymbol{z};\boldsymbol{w}_{GAN})$ is the GAN’s generator, a fully convolutional neural network parameterized by weight vector $\boldsymbol{w}_{GAN}$ that up-convolve the random vector of latent variable $\boldsymbol{z}\sim p_{\boldsymbol{z}}(\boldsymbol{z})$ where $p_{\boldsymbol{z}}$ is a Gaussian zero mean distribution, scaling its dimension up at each layer, to produce microstructure image. A key feature of (4) is that the generator is conditioned on the synthesis inputs $\boldsymbol{\xi}$ as an additional input. During training, $\boldsymbol{z}$ varies randomly, while $\boldsymbol{\xi}$ are set to the corresponding synthesis input used to generate the training data, which consist of microstructural images obtained from FSLBM simulations. Once trainined, $\mathcal{G}_{GAN}(\boldsymbol{\xi},\boldsymbol{z};\boldsymbol{w}_{GAN})$ can efficiently generate random microstructure patterns $\varphi(\mathbf{x})$ by sampling $\boldsymbol{z}$ while by varying $\boldsymbol{\xi}$, the model can produce a wide range of diverse and novel morphologies that correspond to different aerogel synthesis conditions, including those interpolated between training data points.
Training vanilla GANs is notoriously challenging, often requiring significant manual intervention to stabilize the training process. To address these challenges, we utilize Wasserstein GANs with Gradient Penalty (WGAN-GP), as proposed by [42]. WGAN-GP enhances stability by imposing Lipschitz continuity constraints on the discriminator and introducing a gradient penalty term into the discriminator’s loss function. Like vanilla GANs, WGAN-GP is formulated as a two-player zero-sum game, expressed as:
[…]
The loss functions for the generator and discriminator in the WGAN-GP are defined as,
$\mathcal{L}_{g}$ $=$ $-\mathbb{E}_{\boldsymbol{z}\sim p_{z}(\boldsymbol{z})}\left[\mathcal{D}(\mathcal{G}(\boldsymbol{\xi},\boldsymbol{z};\boldsymbol{w}_{GAN});\boldsymbol{w}_{dis})\right]$ (6)
$\mathcal{L}_{d}$ $=$ $\mathbb{E}_{\boldsymbol{z}\sim p_{z}(\boldsymbol{z})}\left[\mathcal{D}(\mathcal{G}(\boldsymbol{\xi},\boldsymbol{z};\boldsymbol{w}_{GAN});\boldsymbol{w}_{dis})\right]-\mathbb{E}_{\varphi(\mathbf{x})\sim p_{data}}\left[\mathcal{D}(\varphi(\mathbf{x});\boldsymbol{w}_{dis})\right]$ (7)
$+$ $c_{\lambda}\mathbb{E}_{\hat{\varphi}(\mathbf{x})\sim p_{\hat{\varphi}(\mathbf{x})}}\left[(\|\nabla_{\hat{\varphi}(\mathbf{x})}\mathcal{D}(\hat{\varphi}(\mathbf{x});\boldsymbol{w}_{dis})\|_{2}-1)^{2}\right],$
where $p_{\hat{\varphi}(\mathbf{x})}$ represents the distribution obtained by uniformly sampling points along straight lines between pairs of samples from $p_{data}$ and $p_{G}$. The discriminator’s loss consists of three components: the first two terms calculate the Wasserstein-1 distance, measuring the difference between real and generated distributions, while the third term imposes a gradient penalty to enforce Lipschitz continuity, with $c_{\lambda}$ as the regularization weight. The generator’s loss minimizes the discriminator’s evaluation of the generated samples, encouraging the generator to produce more realistic outputs [43].
3 Physical and surrogate models of aerogel mechanical properties
This section introduces a physics-based model for analyzing stress and strain distribution within the two-phase aerogel material. Additionally, it describes the development of a microstructure-property surrogate model using a Bayesian convolutional neural network (BayesCNN). Coupled with the microstructural images generated by the WGAN-GP model outlined in Section 2.2, the linked-surrogates enables facilitates efficient prediction of the quantity of interest (QoI), specifically the strain energy.
3.1 Stochastic finite element model of aerogel elastic deformation
To accurately simulate the interplay between microstructural features and macroscopic mechanical properties of ceramic aerogels, we employ a finite element (FE) model governed by a stochastic partial differential equation. This model determines the displacement and stress fields based on the microstructural images of the aerogel. Let $\Omega$ represent a bounded domain in $\mathbb{R}^{d}$, where $d$ can be 1, 2, or 3 with Lipschitz boundary, and let the boundary of this domain be denoted by $\partial\Omega$. The goal is to determine the displacement field $\boldsymbol{u}(\mathbf{x})$ at spatial points $\mathbf{x}\in\Omega$. The governing equation is expressed as:
$\nabla\cdot\boldsymbol{T}(\boldsymbol{u})$ $=$ $\boldsymbol{f}(\mathbf{x}),\quad\mathbf{x}\in\Omega,$
$\boldsymbol{T}(\boldsymbol{u})\boldsymbol{n}$ $=$ $\boldsymbol{t}(\mathbf{x}),\quad\mathbf{x}\in\Gamma_{N},$
$\boldsymbol{u}(\mathbf{x})$ $=$ $\boldsymbol{u}^{*},\quad\mathbf{x}\in\Gamma_{D},$ (8)
where, $\boldsymbol{f}$ is a prescribed source term, $\boldsymbol{t}$ is traction, $\Gamma_{N}$ is the subset of the boundary $\partial\Omega$ where the Neumann boundary condition is applied, $\Gamma_{D}=\partial\Omega\setminus\Gamma_{N}$ represents the boundary subject to Dirichlet conditions. The Cauchy stress tensor $\boldsymbol{T}$ is defined as:
$\boldsymbol{T}(\boldsymbol{u})$ $=$ $2\mu_{s}\,\varphi(\boldsymbol{x})\,\boldsymbol{E}(\boldsymbol{u})+\lambda_{s}\varphi(\boldsymbol{x})\,{\rm tr}(\boldsymbol{E}(\boldsymbol{u}))\boldsymbol{I},$ (9)
[…]
Figure 3: Finite element (FE) simulation results for three microstructure indicator functions, showing the magnitude of displacement and Cauchy stress fields. For the FE simulations, young’s modulus for solid $E_{s}=100\times 10^{6}$ Pa, poisson’s ratio $\nu$ = 0.3 was used. A traction force of $10^{6}$ N was applied in the negative y direction (compression).
The prediction QoI is the scalar strain energy, computed as:
$Q=\int_{\Omega}\boldsymbol{T}(\boldsymbol{u}):\boldsymbol{E}(\boldsymbol{u})\;d\mathbf{x}.$ (10)
3.2 Bayesian convolutional neural network surrogate model
To develop a surrogate model linking microstructure to mechanical properties, we employ a Bayesian convolutional neural network (CNN) that maps a microstructure image of ceramic aerogel $\varphi(\mathbf{x})$ to the corresponding strain energy output $Q^{CNN}=\mathcal{G}_{CNN}(\varphi_{i};\boldsymbol{w}_{CNN})$ (see Figure 1). This mapping is parameterized by the weights $\boldsymbol{w}_{CNN}$ of the filters of the convolutional layers and the fully-connected layers of the CNN architecture, which are determined through training. The training dataset, denoted as $\boldsymbol{D}=\{(\varphi_{i},Q^{FE}_{i})\}_{i=1}^{N_{D}}$ consists of $N_{D}$ samples furnished by the stochastic finite element model described in previous section, such that $Q^{FE}_{i}=\mathcal{F}_{FE}(\varphi_{i};\boldsymbol{\theta}_{FE})$. Traditional CNN training rely on maximum likelihood estimation of the weight parameters to minimize the error between the output and the training data. However, these approaches fail to explicitly account for the uncertainty in CNN weight parameters that translate to uncertainty in the output, which becomes critical in scenarios involving small and uncertain datasets. In contrast, a Bayesian framework provides a more robust approach by inferring probability distribution of the weights from data using the Bayes’ rule,
$\pi_{\text{post }}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})=\frac{\pi_{\text{like }}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})\pi_{\text{prior }}(\boldsymbol{w}_{CNN})}{\pi_{\text{evid }}(\boldsymbol{D})},$ (11)
where $\pi_{\text{prior }}(\boldsymbol{w}_{CNN})$ is prior distribution of the weights, $\pi_{\text{like }}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})$ is the likelihood of observing the data given the weights, and $\pi_{\text{evid }}(\boldsymbol{D})$ is the evidence, a normalization factor ensuring that the posterior integrates to 1. The posterior distribution $\pi_{\text{post }}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})$ reflects the updated knowledge of the weights after incorporating the training data. Given this posterior distribution, for any new microstructure $\varphi^{\star}(\mathbf{x})$ the strain energy output can be predicted using the predictive distribution,
[…]
In practice, a collection of $M_{mc}$ samples are drawn from posterior $\hat{\boldsymbol{w}}_{CNN}^{(j)}\sim\pi_{\text{post }}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})$ and the prediction distribution can then be approximated using Monte Carlo estimation, which is expressed as:
$\pi\left(Q^{CNN\star}\mid\varphi^{\star},\boldsymbol{D}\right)$ $=$ $\mathbb{E}_{\boldsymbol{w}_{CNN}\mid\boldsymbol{D}}\left[\pi\left(Q^{CNN\star}\mid\varphi^{\star},\boldsymbol{w}_{CNN}\right)\right]$ (13)
$\approx$ $\frac{1}{M_{mc}}\sum_{j=1}^{M_{mc}}\mathcal{G}_{CNN}(\varphi^{\star};\hat{\boldsymbol{w}}_{CNN}^{(j)}).$
Variational Inference
To achieve computationally efficient Bayesian inference for high-dimensional parameter spaces, we employ variational inference, an optimization-based technique that approximates the true posterior distribution $\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})$ with a tractable variational distribution $q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$, parameterized by $\boldsymbol{\eta}$ and often chosen from a family of distributions such as Gaussians. The optimal variational distribution $q^{\text{opt}}$ is obtained by minimizing the Kullback-Leibler (KL) divergence between the variational distribution and the true posterior,
[…]
where the KL divergence is defined as:
$\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})\right]$ $=$ $\int q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\ln\frac{q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})}{\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})}d\boldsymbol{w}_{CNN}$ (15)
$=$ $\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\frac{q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})}{\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})}\right].$
Then, the prediction in (12) is performed with replacing the posterior with the tractable variational approximation $q^{\text{opt}}$. Substituting the posterior from Bayes’ theorem (11) into the (15) gives,
$\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})\right]=$displaystyle-- $\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\frac{\pi_{\text{prior}}(\boldsymbol{w}_{CNN})}{q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})}\right]$ (16)
-- $\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})\right]$
$+$ $\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\pi_{\text{evid}}(\boldsymbol{D})\right].$
The evidence $\pi_{\text{evid}}(\boldsymbol{D})$ is independent of the variational distribution $q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$, so the last term in the above equation $\mathbb{E}_{q_{\boldsymbol{\eta}}}[\ln\pi_{\text{evid}}(\boldsymbol{D})]=\ln\pi_{\text{evid}}(\boldsymbol{D})$ and does not contribute to the minimization problem in (14). This leads to the definition of the evidence lower bound (ELBO),
$\mathrm{L}[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]$ $=$ $\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\frac{\pi_{\text{prior}}(\boldsymbol{w}_{CNN})}{q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})}\right]+\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})\right]$ (17)
$=$ $-\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{prior}}(\boldsymbol{w}_{CNN})\right]+\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})\right].$
Thus, the KL divergence can be rewritten as:
$\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{post}}(\boldsymbol{w}_{CNN}\mid\boldsymbol{D})\right]=-\mathrm{L}[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]+\ln\pi_{\text{evid}}(\boldsymbol{D}),$ (18)
where minimizing the KL divergence is equivalent to maximizing $L[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]$ over $q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ and ELBO provides a lower bound on $\ln\pi_{\text{evid}}(\boldsymbol{D})$. Considering the additive noise model, assuming Gaussian-distributed total error resulting from data noise and model inadequacy, the log-likelihood function is given by:
$\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})$ $=$ $-\frac{N_{D}}{2}\ln(2\pi\sigma_{noise})-\frac{1}{2\sigma_{noise}}\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}.$ (19)
where $R(\boldsymbol{w}_{CNN})=\sum_{i=1}^{N_{D}}\left(Q^{CNN}_{i}-Q^{FE}_{i}\right)^{2}$ is called the data misfit and $\sigma_{noise}$ is the noise variance, a hyperparameter that balances the influence of the prior (regularization term) and likelihood (data misfit term) in computing the posterior distribution. The expectation of the log-likelihood is evaluated using Monte Carlo sampling,
$\mathbb{E}_{q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})}[\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})]=\frac{1}{M_{mc}}\sum_{j=1}^{M_{mc}}\ln\pi_{\text{like}}\left(\boldsymbol{D}\mid\hat{\boldsymbol{w}}_{CNN}^{(j)}\right),$ (20)
where $\hat{\boldsymbol{w}}_{CNN}^{(j)}\sim q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})$ are samples drawn from the variational distribution, and $M_{mc}$ is the number of samples. In practice, the variational distribution is expressed as a product of independent Gaussian distributions following the mean-field approximation,
$q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})=\prod_{i=1}^{s}\mathcal{N}(w_{CNN_{i}};\mu_{i},\sigma_{i}),$ (21)
where the variational parameters $\boldsymbol{\eta}$ consist of the means and standard deviations of each Gaussian component, $\boldsymbol{\eta}=\{\mu_{i},\sigma_{i}\}_{i=1}^{d}$. Bayesian inference is then framed as optimizing the variational parameters $\boldsymbol{\eta}$ to ensure that the variational density closely approximates the true posterior distribution. Using the ELBO from (23), this can be written as,
$\boldsymbol{\eta}^{\text{opt }}=\underset{\boldsymbol{\eta}}{\arg\min}\;\;\left\{\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{prior}}(\boldsymbol{w}_{CNN})\right]-\mathbb{E}_{q_{\boldsymbol{\eta}}}\left[\ln\pi_{\text{like}}(\boldsymbol{D}\mid\boldsymbol{w}_{CNN})\right]\right\}.$ (22)
Considering Gaussian prior, the first term in (22) has a closed-form solution, while the second term can be evaluated using (20). For numerical optimization, the ELBO from (17) is reformulated as,
$\mathrm{L}[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})]\propto\beta\mathrm{KL}\left[q_{\boldsymbol{\eta}}(\boldsymbol{w}_{CNN})\|\pi_{\text{prior}}(\boldsymbol{w}_{CNN})\right]+\frac{1}{M_{mc}}\sum_{j=1}^{M_{mc}}R(\hat{\boldsymbol{w}}_{CNN}^{(j)}),$ (23)
where $\beta=2\sigma_{\text{noise}}$ and the right-hand side of the (23) is employed in the optimization process to determine the variational parameters. To enhance the efficiency of gradient-based optimization, we applied the local reparameterization trick to convolutional layers. This approach preserves computational performance while enabling efficient updates of variational parameters. The optimization technique, commonly referred to as “Bayes by Backprop” [17], facilitates the training of BayesCNN by leveraging this reparameterization framework.
4 Numerical results
This section presents computational results from the integrated modeling framework (Figure 1), focusing on the development and evaluation of generative and BayesCNN surrogate models. Their predictive accuracy is assessed against high-fidelity methods, including the Free-surface Lattice Boltzmann Method (FSLBM) and stochastic Finite Element (FE) simulations. The linked surrogates are then used to propagate uncertainty through the synthesis-microstructure-property relationships, analyzing the impact of microstructural randomness and surrogate model errors on the reliability of the predicted quantities of interest (QoI).
4.1 Generative surrogate model of microstructure
We detail the construction of the WGAN-GP surrogate model and quantitatively evaluate its performance by comparing microstructure images generated by the WGAN-GP to the synthetic images from FSLBM simulations.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Networks architecture and training
The WGAN-GP generator network consists of four layers, each with a Convolutional Transpose, Batch Normalization, and ReLU activation, followed by a final layer with a Convolutional Transpose and Tanh activation to output the generated microstructure image. A latent vector $\boldsymbol{z}$ of size 4 $\times$ 1 is progressively upscaled by a factor of 2 in each subsequent layer to produce the 128 $\times$ 128 $\times$ 1 microstructure. The discriminator network mirrors this structure, using Convolutional, Batch Normalization, and Leaky ReLU layers, with a final Convolutional layer producing a scalar output representing the Wasserstein distance between synthetic and generated image distributions. Both networks employ a 4 $\times$ 4 kernel, 2 $\times$ 2 stride, and 1 $\times$ 1 padding, achieving a scaling factor of $2^{5}$ (=32) over five layers. The architectures, summarized in Table 1 (see Appendix) , are adapted from previous GAN-based material microstructure models [37, 38], with modifications tailored to this study. The training dataset comprises 128 $\times$ 128 pixel microstructure images created using FSLBM simulations, $\mathcal{F}_{LB}(\boldsymbol{\xi};\boldsymbol{\theta}_{BL})$. The WGAN-GP was trained on 2,500 synthetic images from FSLBM simulations (Figure 2(a)) with an average pore size of $6\mu m$. Training employed the Adam optimizer with a learning rate of 0.0002, a batch size of 10, and a discriminator penalty coefficient $c_{\lambda}=10$, over 2000 epochs.
Accuracy of the generated microstructure images
Figure 4 illustrates a comparison between synthetic microstructure images of aerogels from FSLBM simulations and those generated by the WGAN-GP model. The accuracy of the WGAN-GP surrogate model is quantitatively assessed by comparing morphological attributes of generated microstructures with synthetic ones from FSLBM simulations, focusing on pore size distribution and the two-point correlation function. Figure 5(a) shows the Kernel Density Estimate (KDE) of pore diameters for 50 samples, revealing close agreement between synthetic and generated microstructures. Pore diameters were calculated using an image segmentation method that labels individual pores, determines their areas (based on pixel counts), and computes the equivalent diameter of a circle with the same area. Further details on this method are provided in [45].
(a)                                              (b)
Figure 5: Morphological features of 50 samples from synthetic and generated microstructure images: (a) Probability distribution of pore diameters. (b) 1D plots of the two-point correlation function.
The two-point correlation function, $S_{2}(\boldsymbol{r})$, is a statistical measure characterizing the morphology of two-phase media [46], representing the probability that two points separated by $\lvert\boldsymbol{r}\rvert$ belong to the same phase. Here, we focus on the pore phase (pixels with a value of 255) to quantify pore distribution within a given distance $\boldsymbol{r}$. Figure 5(b) shows 1D plots of $S_{2}(\boldsymbol{r})$ for 50 synthetic and generated samples, with error bars capturing sample variability. The close agreement demonstrates the WGAN-GP’s ability to reproduce microstructural features consistent with physical simulations.
Scalable microstructure generator
Despite the widespread application of GANs in modeling material microstructures, their utility is fundamentally constrained by the fixed size of training images. However, material properties of practical interest are often required at the component level, involving domains several orders of magnitude larger than those captured by microscale imaging or modeling techniques. To overcome this limitation, we propose a novel approach that enables the generation of arbitrary-sized microstructure images using WGAN-GP trained on smaller image datasets. This method utilizes a latent vector as input to the generator, trained in a low-dimensional latent space, allowing efficient generation of microstructures at various scales while maintaining control over the output dimensions.
(a) 96$\times$96             (b) 128$\times$128             (c) 160$\times$160
Refer to caption Refer to caption Refer to caption
(d) 192$\times$192             (e) 384$\times$384             (f) 1280$\times$1280
Figure 6: Generated microstructural images of varying dimensions (in pixels) produced by the scalable generator network, corresponding to ceramic aerogel domain sizes of (a) 52.5 $\mu m$, (b) 70 $\mu m$, (c) 87.5 $\mu m$, (d) 105 $\mu m$, (e) 210 $\mu m$, and (f) 700 $\mu m$.
As described, the generator architecture, trained with a latent vector $\boldsymbol{z}$ of size 4 $\times$ 1, provides an upscaling factor of 2 per layer. However, increasing the image dimensions requires adding more convolutional layers, which also necessitates corresponding layers in the discriminator. This increases network complexity, training time, and computational cost. To address these issues, we reformulate the generator architecture by separating it from the fully connected layers and using linear transformations to reshape the latent vector $\boldsymbol{z}$. Specifically, $\boldsymbol{z}$ of dimension 4 $\times$ 1 is mapped to a latent variable of dimension (-1, hidden_dim, 4, 4) where hidden_dim is the hidden dimension size, and -1 denotes the arbitrary batch size. This approach provides more flexibility over the size of the latent vector we input to the generator, and this also results in the reduction of the total number of layers. Importantly, this method preserves the microstructural features while allowing for the efficient generation of images at arbitrary dimensions. Figure 6 shows microstructures of various sizes generated by the scalable WGAN-GP, trained on 128$\times$128 pixel synthetic microstructures (corresponding to a 70 $\mu m$ aerogel domain). The output size is controlled by the dimensions of the latent variable, providing flexibility. Figure 7 compares the KDEs of pore sizes from 30 synthetic microstructures and larger generated images. The agreement highlights the scalable generator’s ability to produce larger microstructure images while preserving key statistical features across different sizes.
(a)                                                         (b)
Figure 7: Comparison of pore diameter probability distributions between synthetic 128 $\times$ 128 microstructures and scaled generated microstructures with dimensions: (a) 640 $\times$ 640 pixels (350 $\mu m$),  (b) 1280 $\times$ 1280 pixels (700 $\mu m$).
4.2 BayesCNN surrogate model of mechanical properties
This section describes the construction of the BayesCNN surrogate model, using high-fidelity simulations from the stochastic FE model, to predict the mapping between microstructure images and the strain energy of ceramic aerogels (QoI). The surrogate model’s performance with quantified uncertainty, is validated for both in-distribution and out-of-distribution predictions. In-distribution refers to microstructures with pore sizes and morphology similar to the training data, while out-of-distribution includes microstructures with characteristics that interpolate between the training data.
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: High-fidelity training data to construct the BayesCNN surrogate model. Probability distribution of strain energy $Q^{FE}$ computed using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ and samples of three microstructure categories: small pores with average pore sizes of $6\mu m$; medium pores with average pore sizes of $18\mu m$; large pores with average pore sizes of $25\mu m$.
High-fidelity data
Three categories of ceramic aerogel microstructures: small pores, medium pores, and large pores, with average pore sizes of $6\mu m$, $18\mu m$, and $25\mu m$, respectively, is taken into account (Figure 8). A total of 1,188 samples (396 per category) were simulated using the stochastic FE model $\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ to compute the strain energy $Q^{FE}$. For all samples, uniaxial loading conditions are applied with traction of ${t}_{y}=-10^{6}N$ being imposed downward on the top surface, while the displacement at the bottom boundary of the domain are fixed. The resulting microstructure-strain energy data was randomly divided into 90% training dataset and 10% in-distribution dataset.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Finite element results for small pores microstructures: (a) typical high-strain-energy sample (98% of cases), and (b) anomalous low-strain-energy sample, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Finite element results for medium pores microstructures: (a) typical high-strain-energy sample (97% of cases), and (b) anomalous low-strain-energy sample, showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Finite element results for large pores microstructures: (a) low-strain-energy sample (25% of cases), and (b) medium-strain-energy sample (25% of cases), and (c) high-strain-energy sample (25% of cases), showing the microstructure indicator function $\phi(\mathbf{x})$, the magnitude of displacement $\|\mathbf{u}\|$, and components of Cauchy stress $\mathbf{T}$.
Additionally, to assess the BayesCNN’s generlizability, an out-of-distribution dataset was generated using FSLBM simulations. This dataset includes 10 microstructures with an average pore size of 12 $\mu m$ (intermediate between small pores and medium pores) and 10 with an average pore size of 22 $\mu m$ (between medium pores and large pores). The FSLBM parameters matched those used for training data generation, except for the average pore size and initial bubble count: 75 bubbles for the 12 $\mu m$ microstructures and 35 for the 22 $\mu m$ cases.
The bimodal distribution of strain energy probabilities, illustrated in Figure 8(d), reflects the distinct mechanical behaviors exhibited within and across the three categories of ceramic aerogel microstructures. Representative FE solutions for these categories are presented in Figures 911. For the small pores category, the majority of microstructural samples (approximately 98%) exhibit strain energies ranging from $1000\times 10^{-7}$ to $1100\times 10^{-7}\,\text{mJ}$ with occasional anomalies at lower strain energy levels. As shown in Figure 9, most microstructures demonstrate mechanical deformation dominated by high-strain pathways connecting the top and bottom boundaries. In some cases, these pathways are interrupted by isolated solid-phase regions, leading to deviations in mechanical response. A similar trend is observed for the medium pores category, where strain energies predominantly range from $1150\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$, with sporadic anomalies indicating increased mechanical strength (Figure 10). In contrast, the large pores category exhibits a distinct mechanical behavior characterized by a near-normal strain energy distribution ranging from $800\times 10^{-7}$ to $1300\times 10^{-7}\,\text{mJ}$ (Figure 11). These variations highlight the strong influence of pore size and distribution on the mechanical properties of ceramic aerogels.
Network architecture and training
The BayesCNN’s architecture and hyperparameters were determined through a systematic validation process using the in-distribution dataset (10% of high-fidelity data generated using stochastic FE model). Table 2 (see Appendix) details the architecture, including input and output shapes and the total number of parameters. Key hyperparameters for the variational inference include the learning rate employed by the Adam optimizer and $\beta$ in (23), which balances data misfit and prior regularization in the Bayesian framework.
Figure 12 illustrates the process of selecting these hyperparameters. The validation loss, defined as the mean squared error between validation data and the corresponding BayesCNN mean predictions, is plotted over training epochs for four different learning rates in Figure 12(a). The four learning rates yield similar validation losses over the chosen course of epochs. A learning rate of 0.002, yielding the most stable validation loss decrease (fewest number of spikes), was chosen for Bayesian training. Additionally, the validation loss stopped decreasing around 400 epochs, corresponding to the minimum loss. The accuracy of Bayesian inference depends critically on the proper selection of $\beta$. Overestimating $\beta$ diminishes the influence of valuable data, while underestimating it leads to over-fitting, biased parameter estimates. To avoid such biases, $\beta$ was tuned based on its impact on $R^{2}$ values computed as,
[…]
where $N_{id}$ is the number of in-distribution datapoints, $Q^{FE}$ is the deterministic strain energy computed from FE model for each microstructure sample, and $Q^{CNN}$ the BayesCNN prediction computed using 40 samples of the weight parameters for each microstructure. As shown in Figure 12(b), $\beta=0.01$ consistently achieves the highest $R^{2}$ for both in-distribution and out-of-distribution data, and thus selected for the Bayesian inference. This result also aligns with the cold posterior effect [47], demonstrating improved predictive performance for $\beta$ below 1.
(a)                                                         (b)
Figure 12: Hyperparameter selection for Bayesian inference: (a) Validation loss as a function of training epochs for various learning rates used in the optimizer. (b) Comparison of $R^{2}$ values across different $\beta$ values in both in-distribution and out-of-distribution datasets.
[…]
BayesCNN prediction
Figures 1315 show the in-distribution prediction results of the BayesCNN model for the small pores, medium pores, and large pores microstructure classes. Each figure includes the probability distribution of high-fidelity strain energy data for the corresponding class and the three representative microstructure samples including the true strain energy value from the FE model, $Q^{FE}=\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$ and the BayesCNN-predicted probability distribution, $Q^{CNN}=\mathcal{G}_{CNN}(\varphi,\boldsymbol{w}_{CNN})$, obtained using 40 samples of weight parameters $\boldsymbol{w}_{CNN}$. In all cases, the true values consistently fall within the BayesCNN-predicted distributions, demonstrating the model’s reliability for in-distribution predictions.
[…]
A key advantage of deep learning surrogate models for microstructure-property mapping in materials design and engineering lies in their accuracy and reliability for out-of-distribution predictions. This refers to the model’s ability to generalize to new microstructural features beyond the classes represented in the training data. Figure 16 compares the true strain energy, $Q^{FE}=\mathcal{F}_{FE}(\varphi,\boldsymbol{\theta}_{FE})$, computed via the FE model, with the probability distribution of strain energy predicted by the BayesCNN surrogate model, $Q^{CNN}=\mathcal{G}_{CNN}(\varphi,\boldsymbol{w}_{CNN})$, for a few samples.
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to caption   Refer to caption
(b)                                                         (c)
Figure 16: Out-of-distribution predictions of the BayesCNN model: (a) Microstructure samples with average pore sizes of 12 $\mu m$ (between small pores and medium pores) and 22 $\mu m$ (between medium pores and large pores), along with the strain energy probability distribution of out-of-distribution dataset. (b) BayesCNN predictions for two microstructures with 12 $\mu m$ pores. (c) BayesCNN predictions for two microstructures with 2 $\mu m$ pores.
The accuracy and reliability of the BayesCNN predictions are further evaluated for both in-distribution and out-of-distribution dataset in Figure 17. In this figure, the prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is represented by the variance of the BayesCNN predictions, $\mathbb{V}[Q^{CNN}]$. These results indicate that the BayesCNN model successfully identifies out-of-distribution data, as reflected by lower accuracy and reliability in its predictions for such cases. The model performs well in interpolating between small pores and medium pores microstructures. However, its accuracy diminishes for out-of-distribution data with an average pore size of 22 $\mu m$, representing microstructures between medium pores and large pores.
Figure 17: The BayesCNN model’s prediction accuracy and reliability for in-distribution and out-of-distribution datasets using 10 samples. The prediction error is defined as $\left|\mathbb{E}[Q^{CNN}]-Q^{FE}\right|$, and the uncertainty is defined as $\mathbb{V}[Q^{CNN}]$.
5 Conclusions
This study introduces an integrated deep learning surrogate modeling framework for efficient uncertainty propagation across the synthesis, microstructure, and properties of ceramic aerogel porous materials. The stochastic microstructure surrogate employs a scalable Wasserstein Generative Adversarial Network with Gradient Penalty (WGAN-GP) to generate ensembles of morphologically consistent microstructures over arbitrarily large domains, trained on smaller domain size samples from Lattice Boltzmann simulations of the foaming process. For property prediction, a Bayesian Convolutional Neural Network (BayesCNN) is developed to estimate weight parameter distributions using small, uncertain high-fidelity datasets from finite element elasticity simulations. Numerical results demonstrate that the WGAN-GP effectively generates microstructures consistent with the morphology of training data (in terms of pore size distribution and two-point correlation) across larger domains. The BayesCNN accurately maps microstructural images to strain energy while quantifying uncertainty, achieving reliable predictions for unseen in-distribution data. However, its predictive accuracy and confidence decline for out-of-distribution microstructures with morphologically distinct attributes. Notably, while the model performs well in interpolating between small and medium pore sizes, it struggles with medium-to-large pore sizes due to significant differences in the deformation mechanisms of aerogels with larger pores.
Future work will focus on enhancing the framework by incorporating aerogel synthesis parameters as additional inputs to the WGAN-GP, enabling the discovery of novel morphologies that interpolate between training data for specific synthesis parameter combinations, as demonstrated in [37]. The BayesCNN will be further evaluated across a wider range of aerogel microstructures, investigating both interpolation and extrapolation capabilities. This includes assessing predictive accuracy for microstructures with diverse pore size distributions and integrating nonlinear finite element models based on continuum damage mechanics to capture complex deformation behaviors of ceramic aerogels. Additionally, the surrogate model will be refined through improved architecture design and hyperparameter selection using a hierarchical Bayesian inference framework [18]. This approach will identify the most plausible BayesCNN model under constraints of limited and uncertain training data, enhancing the framework reliable deployment in high-consequence material discovery scenarios.
This study tackles the computational challenges of model-driven materials discovery by developing a validated deep learning surrogate model for synthesis-microstructure-property relationships using limited and uncertain high-fidelity data. It also paves the way for advancing predictive accuracy and confidence in data-driven surrogate modeling for physical and engineering systems.
Acknowledgments
MAI and DF extend their gratitude for the financial support of this work from the Research Foundation of The State University of New York (SUNY) under grant number 1191358. Additionally, the authors would like to acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
Appendix A Appendix: Deep learning architectures
Layer Dimension Kernel Stride Padding Batch Normalization Activation Function
Generator
Layer 1 -1 $\times$ 8 $\times$ 8 $\times$ 512 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes ReLU
Layer 2 -1 $\times$ 16 $\times$ 16 $\times$ 256 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes ReLU
Layer 3 -1 $\times$ 32 $\times$ 32 $\times$ 128 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes ReLU
Layer 4 -1 $\times$ 64 $\times$ 64 $\times$ 64 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes ReLU
Layer 5 -1 $\times$ 128 $\times$ 128 $\times$ 1 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 No Tanh
Discriminator
Layer 1 -1 $\times$ 64 $\times$ 64 $\times$ 64 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes LeakyReLU
Layer 2 -1 $\times$ 32 $\times$ 32 $\times$ 128 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes LeakyReLU
Layer 3 -1 $\times$ 16 $\times$ 16 $\times$ 256 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes LeakyReLU
Layer 4 -1 $\times$ 8 $\times$ 8 $\times$ 512 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 Yes LeakyReLU
Layer 5 -1 $\times$ 1 $\times$ 1 $\times$ 1 4 $\times$ 4 2 $\times$ 2 1 $\times$ 1 No -
Table 1: Architecture of WGAN-GP with the output dimension from each layer.
Layer Type Kernel Stride Input Shape Output Shape Parameters
Input - - - $3\times 168\times 168$ $3\times 168\times 168$ 0
Conv1 BBBConv2d $5\times 5$ $1\times 1$ $3\times 168\times 168$ $6\times 164\times 164$ $456$
Act1 Softplus - - $6\times 164\times 164$ $6\times 164\times 164$ 0
Pool1 MaxPool $2\times 2$ $2\times 2$ $6\times 164\times 164$ $6\times 82\times 82$ 0
Conv2 BBBConv2d $5\times 5$ $1\times 1$ $6\times 82\times 82$ $16\times 78\times 78$ $2{,}416$
Act2 Softplus - - $16\times 78\times 78$ $16\times 78\times 78$ 0
Pool2 MaxPool $2\times 2$ $2\times 2$ $16\times 78\times 78$ $16\times 39\times 39$ 0
Flatten Flatten - - $16\times 39\times 39$ $24{,}336$ 0
FC1 BBBLinear - - $24{,}336$ $120$ $2{,}920{,}440$
Act3 Softplus - - $120$ $120$ 0
FC2 BBBLinear - - $120$ $84$ $10{,}164$
Act4 Softplus - - $84$ $84$ 0
Regression BBBLinear - - $84$ $1$ $85$
Table 2: Architecture of the BayesCNN model and total number of parameters.